# Understanding the Cluster LP for Correlation Clustering

### v2

@@ 4.4 Construction of Solution to the Cluster LP Using Independently Sampled Clusters
For every $u\in V$, let $R_{u}=\{t:u\in C_{t}\}$. Notice that $\Delta y_{\emptyset}\cdot\frac{|E_{\mathrm{adm}}|}{y_{\emptyset}n^{2}}=\Theta\big{(}\frac{\log n}{\varepsilon_{1}^{2}}\big{)}$, with a large enough hidden constant. Using Chernoff bound and union bound, we can prove that with probability at least $1-1/n$, the following conditions hold.
@@
In the second stage we do the following. While $|R_{u}|=|R_{v}|>{\left\lceil(1-\varepsilon)\Delta\right\rceil}$, we remove the largest index in $R_{u}$ from $R_{u}$, and do the same for $R_{v}$. Consider one iteration of the while loop. If the two indices are the same, then $|R_{u}\setminus R_{v}|=|R_{v}\setminus R_{u}|$ does not change. Otherwise, wlog we assume the index we removed from $R_{u}$ is larger. Then removing the index in $R_{u}$ will decrease $|R_{u}\setminus R_{v}|$. So the iteration can not increase $|R_{u}\setminus R_{v}|=|R_{v}\setminus R_{u}|$. ∎
@@
$\mathrm{obj}(\tilde{x})$ $\leq(1+O(\varepsilon))\left(\mathrm{obj}(x)+y_{\emptyset}\sum_{uv\in{V\choose 2}}\mathrm{err}_{uv}\right)+O(\varepsilon_{1})|E_{\mathrm{adm}}|\leq(1+O(\varepsilon))\mathrm{obj}(x)+O(\varepsilon_{1})|E_{\mathrm{adm}}|$
@@
5 Triangle Analysis used in 1.5-Approximation: Proof of Lemma 9
The goal of this section is to prove Lemma 9, which is restated here. See 9 The proof of Lemma 9 follows from Lemma 25, 26, 27, 28, 29 which we prove in this section. First, we recall some notation.
Notations and Useful Observations
We define some notation and make some observations that will be useful in this section and in Sections 6 and 7. Recall that we are given a polynomial-sized LP solution $(z=(z_{S})_{S\subseteq V},x\in[0,1]^{V\choose 2})$ to (cluster LP). For a $+$edge $uv$, we say $uv$ is short if $x_{uv}\leq\frac{1}{3}$ and long if $x_{uv}>\frac{1}{3}$. Notice that the pivot-based rounding procedure (Algorithm 1) treat short and long $+$edges differently. For vertices $u,v,w\in V$, we define
@@
$y_{u|v|w}$ $:=1-y_{uv|w}-y_{uw|v}-y_{vw|u}-y_{uvw}=1-y_{uv}-y_{uw}-y_{vw}+2y_{uvw}\in[0,1].$
To see that $y_{u|v|w}\in[0,1]$, notice that $y_{uv}+y_{uw}+y_{vw}-2y_{uvw}=\sum_{S}z_{S}\cdot(1_{\{u,v\}\subseteq S}+1_{\{u,w\}\subseteq S}+1_{\{v,w\}\subseteq S}-2\cdot 1_{\{u,v,w\}\subseteq S})$. The quantity after $z_{S}$ is either 00 or $1$, depending on the size of $S\cap\{u,v,w\}$. In an integral solution $z$, $y_{uv},y_{uvw},y_{uv|w}$ and $y_{u|v|w}$, respectively, indicate if $u$ and $v$ are in the same cluster, if $u,v$ and $w$ are all in the same cluster, if $u$ and $v$ are in the same cluster not containing $w$, and if $u,v,w$ are in three different clusters. These definitions also hold when $u,v$ and $w$ are not distinct vertices.
@@
$\Pr[\{v,w\}\subseteq C\mid u]$ =$=\begin{cases}y_{uvw}&\text{if both $uv$ and $uw$ are long $+$edges}\\ \Pr[v\in C\mid u]\cdot\Pr[w\in C\mid u]&\text{otherwise}\end{cases},$
$\Pr[v\in C\wedge w\notin C\mid u]$ =$=\begin{cases}y_{uv|w}&\text{if both $uv$ and $uw$ are long $+$edges}\\ \Pr[v\in C\mid u]\cdot\big{(}1-\Pr[w\in C\mid u]\big{)}&\text{otherwise}\end{cases}.$
We now prove some useful facts about the budget functions defined in Section 2,
$b^{+}_{\alpha}(x)=\frac{\alpha}{1-\alpha/2}\cdot\frac{x^{2}}{1+x},\qquad\text{and}\qquad b^{-}_{\alpha}(x)=\frac{\alpha}{1-\alpha/2}\cdot\frac{(1+2x)(1-x)}{2(1+x)},\qquad\forall x\in[0,1].$
@@
For any $\alpha\in[1,2)$, $b^{+}_{\alpha}(x)$ is convex and decreasing. Moreover, $b^{-}_{\alpha}(x)$ is increasing.
@@
Consider the first and second derivative of $b^{+}_{\alpha}(x)$,
$\frac{d}{dx}b^{+}_{\alpha}(x)=\frac{\alpha}{1-\alpha/2}\cdot\frac{x(2+x)}{(1+x)^{2}}\geq 0,$
$\frac{d^{2}}{d^{2}x}b^{+}_{\alpha}(x)=\frac{\alpha}{1-\alpha/2}\cdot\frac{2}{(x+1)^{3}}\geq 0.$
Here, we used that $\frac{\alpha}{1-\alpha/2}\geq 0$ and $x\geq 0$. Similar, for $b^{-}_{\alpha}(x)$,
$\frac{d}{dx}b^{-}_{\alpha}(x)=\frac{\alpha}{1-\alpha/2}\cdot\frac{-x(x+2)}{(1+x)^{2}}\leq 0.$
Now we start the proof of Lemma 9. Note that for $\alpha=1.5$, the budget functions simplify to
$b^{+}(x):=\frac{6x^{2}}{1+x},\qquad\text{and}\qquad b^{-}(x):=\frac{3(1+2x)(1-x)}{(1+x)},\qquad\forall x\in[0,1].$
We will prove the lemma by considering each possible triangle type separately. That is, degenerate triangles, triangles with no $+$edge (-\!-\!-- - - triangles), triangles with one $+$edge (++\!-\!-+ - - triangles), triangles with two $+$edges ($+$\!+\!-+ + - triangles) and lastly, triangles with three $+$edges (++++\!+\!++ + + triangles). Moreover, recall that the pivot rounding algorithm treats short $+$edges differently than long $+$edges. Therefore, for each type of triangle, we consider different cases that specify how many short and long $+$edges the triangle contains.
Degenerate triangles
@@
Fix budget functions $b^{+}\equiv b^{+}_{\alpha}$ and $b^{-}\equiv b^{-}_{\alpha}$, where $\alpha\geq 4/3$. For any degenerate triangle $T$, we have $\mathrm{cost}(T)\leq\Delta(T)$.
@@
For a --edge with value $x$ this is equivalent to
@@
This holds true since $\frac{\alpha(1+2x)}{(2-\alpha)(1+x)}\geq\frac{\alpha}{(2-\alpha)}\geq 1$, for any $\alpha\geq 1$. For short $+$edges the cost is 0 and the inequality holds trivially. For long $+$edges with value $x>\frac{1}{3}$ we need,
@@
This holds true since $\frac{2\alpha x^{2}}{(2-\alpha)(1+x)}\geq\frac{2\alpha/3}{(2-\alpha)(1+1/3)}\geq 1$ for any $\alpha\geq 4/3$. ∎
5.1 ++++\!+\!++ + + Triangles
@@
Fix the budget function $b^{+}\equiv b^{+}_{\alpha}$, where $\alpha=1.5$. For any ++++\!+\!++ + + triangle $T$, we have $\mathrm{cost}(T)\leq\Delta(T)$.
@@
We distinguish between the following cases: the triangle has three short edges, two short edges, one short edge or no short edge.
(s,s,s)
@@
For a ++++\!+\!++ + + triangle $T=(u,v,w)$, assume that $uv$ is long and $uw$, $vw$ are short:
$\mathrm{cost}(T)$ $=2\cdot x_{uv},$
$\Delta(T)$ $=b^{+}(x_{uv})+b^{+}(x_{uw})+b^{+}(x_{vw})\geq b^{+}(x_{uv})+2\cdot b^{+}\left(\frac{x_{uw}+x_{vw}}{2}\right)\geq b^{+}(x_{uv})+2\cdot b^{+}\left({\frac{x_{uv}}{2}}\right).$
Here, we used that the function $b^{+}$ is convex (Lemma 24) together with the triangle inequality, $x_{uv}\leq x_{uw}+x_{vw}$. For $x_{uv}>1/3$, we have that
$\Delta(T)\geq b^{+}(x_{uv})+2\cdot b^{+}\left({\frac{x_{uv}}{2}}\right)\geq\frac{3}{2}x_{uv}+\frac{7}{6}x_{uv}\geq 2\cdot x_{uv}=\mathrm{cost}(T).$
@@
For a ++++\!+\!++ + + triangle $T=(u,v,w)$, assume that $uv$ is short and $uw$ and $vw$ are long:
$\mathrm{cost}(T)$ $=x_{uw}+x_{vw}+y_{vw}+y_{uw}-2\cdot y_{uvw}$
$\Delta(T)$ $=b^{+}(x_{uw})+b^{+}(x_{vw})+b^{+}(x_{uv})\cdot(y_{vw}+y_{uw}-y_{uvw})$
$\geq b^{+}(x_{uw})+b^{+}(x_{vw}).$
Because $x_{uw}>1/3$,
$b^{+}(x_{uw})\geq\frac{3}{2}x_{uw}=\frac{3}{2}(1-y_{uw}).$
Hence,
$\Delta(T)-\mathrm{cost}(T)$ $\geq\frac{1}{2}(x_{uw}+x_{vw})-y_{vw}-y_{uw}+2\cdot y_{uvw}$
$=1-\frac{3}{2}(y_{vw}+y_{uw})+2\cdot y_{uvw}$
$\geq y_{uv}+y_{vw}+y_{uw}-2\cdot y_{uvw}-\frac{3}{2}(y_{vw}+y_{uw})+2\cdot y_{uvw}$
$=y_{uv}-\frac{1}{2}(y_{vw}+y_{uw})$
@@
For the first inequality we used $1-(y_{uv}+y_{vw}+y_{uw}-2\cdot y_{uvw})\geq 0$ and for the second that $y_{vw},y_{uw}\leq\frac{2}{3}$ and $y_{uv}\geq\frac{2}{3}$.
@@
For a ++++\!+\!++ + + triangle $T$ where we use correlated rounding, [CLN22] show that $\frac{\mathrm{cost}(T)}{\Delta(T)}\leq 1$, when we use budget $\frac{3}{2}x$ for each $+$edge. We have $b^{+}(x)\geq\frac{3}{2}x$ for $x\geq 1/3$. ∎
5.2 -\!-\!-- - - Triangles
@@
Fix the budget function $b^{-}\equiv b^{-}_{\alpha}$, where $\alpha\geq 1$. For any -\!-\!-- - - triangle $T$, we have $\mathrm{cost}(T)\leq\Delta(T)$.
@@
We have that $b^{-}_{\alpha}(x)=\frac{\alpha}{1-\alpha/2}\cdot\frac{(1+2x)(1-x)}{2(1+x)}\geq(1-x)$. As before, $\frac{\alpha(1+2x)}{(2-\alpha)(1+x)}\geq\frac{\alpha}{(2-\alpha)}\geq 1$, for any $\alpha\geq 1$[CLLN23] shows that the inequality holds true if the coefficient is $1$.
@@
Fix budget functions $b^{+}\equiv b^{+}_{\alpha}$ and $b^{-}\equiv b^{-}_{\alpha}$, where $\alpha\geq\frac{4}{3}$. For any ++\!-\!-+ - - triangle $T$, we have $\mathrm{cost}(T)\leq\Delta(T)$.
@@
We have to prove the following two cases.
short $+$edge
For a ++\!-\!-+ - - triangle $T=(u,v,w)$, assume that $uv$ is a short $+$edge and $uw$ and $vw$ are --edges.
$b^{-}_{\alpha}(x)=\frac{\alpha}{1-\alpha/2}\cdot\frac{(1+2x)(1-x)}{2(1+x)}\geq 2(1-x)$
@@
$\Delta(T)$ $\geq b^{-}(x_{uw})+b^{-}(x_{vw})\geq 2(y_{uw}+y_{vw}).$
long $+$edge
For a ++\!-\!-+ - - triangle $T=(u,v,w)$, assume that $uv$ is a long $+$edge and $uw$ and $vw$ are --edges. For $x\geq 1/3$, we can lower bound the budget of the $+$edge,
$b^{+}_{\alpha}(x)=\frac{\alpha}{1-\alpha/2}\cdot\frac{x^{2}}{1+x}\geq\frac{2\alpha/3}{(2-\alpha)(1+1/3)}x\geq x.$
As above, $b^{-}(x)\geq 2(1-x)$. The inequality is true if the coefficient for the $+$edge is $1$ and for the --edge is $2$, [CLLN23].
$\mathrm{cost}(T)$ $=y_{uw}y_{uv}+y_{vw}y_{uv}+y_{uw}+y_{vw}-2y_{uw}y_{vw}=(2-x_{uv})(y_{uw}+y_{vw})-2y_{uw}y_{vw}.$
$\Delta(T)$ $\geq(y_{uw}+y_{vw}-y_{uw}y_{vw})x_{uv}+2(y_{uw}+y_{uv}-y_{uw}y_{uv})y_{vw}+2(y_{vw}+y_{uv}-y_{vw}y_{uv})y_{uw}$
$=(y_{uw}+y_{vw}-y_{uw}y_{vw})x_{uv}+2(1-x_{uv}+y_{uw}x_{uv})y_{vw}+2(1-x_{uv}+y_{vw}x_{uv})y_{uw}$
$=(2-x_{uv})(y_{uw}+y_{vw})+3y_{uw}y_{vw}x_{uv}.$
5.4 $+$\!+\!-+ + - Triangles

### v3

@@ 4.3 Using Independently Sampled Clusters to Construct Cluster LP Solution
For every $u\in V$, let $R_{u}=\{t:u\in C_{t}\}$. Notice that $\Delta y_{\emptyset}\cdot\frac{|E_{\mathrm{adm}}|}{y_{\emptyset}n^{2}}=\Theta\big(\frac{\log n}{\varepsilon_{1}^{2}}\big)$, with a large enough hidden constant. Using Chernoff bound and union bound, we can prove that with probability at least 11/n1-1/n, the following conditions hold.
@@
In the second stage we do the following. While $|R_{u}|=|R_{v}|>{\left\lceil(1-\varepsilon_{1})\Delta\right\rceil}$, we remove the largest index in $R_{u}$ from $R_{u}$, and do the same for $R_{v}$. Consider one iteration of the while loop. If the two indices are the same, then $|R_{u}\setminus R_{v}|=|R_{v}\setminus R_{u}|$ does not change. Otherwise, wlog we assume the index we removed from $R_{u}$ is larger. Then removing the index in $R_{u}$ will decrease $|R_{u}\setminus R_{v}|$. So the iteration can not increase $|R_{u}\setminus R_{v}|=|R_{v}\setminus R_{u}|$. ∎
@@
[…]
@@
5 Triangle Analysis used in 1.56-Approximation of Cluster LP
The goal of this section is to prove Lemma 9, which is restated here.
See 9 The proof of Lemma 9 follows from Lemma 35, 36, 37, 38, 39 which we prove in this section. First, we recall some notation.
Notation and Useful Observations.
We define some notation and make some observations that will be useful in this section and in Section 6. Recall that we are given a polynomial-sized LP solution $(z=(z_{S})_{S\subseteq V},x\in[0,1]^{V\choose 2})$ to (cluster LP). Moreover, remember that we choose the threshold $\tau=0.4$. For a ++edge uvuv, we say $uv$ is short if $x_{uv}\leq\tau$ and long if $x_{uv}>\tau$. Notice that the pivot-based rounding procedure (Algorithm 1) treats short and long ++edges differently. For vertices $u,v,w\in V$, we define
@@
$y_{u|v|w}$ $:=1-y_{uv|w}-y_{uw|v}-y_{vw|u}-y_{uvw}.$
In an integral solution $z$, $y_{uv},y_{uvw},y_{uv|w}$ and $y_{u|v|w}$, respectively, indicate if $u$ and $v$ are in the same cluster, if $u,v$ and $w$ are all in the same cluster, if $u$ and $v$ are in the same cluster not containing $w$, and if $u,v,w$ are in three different clusters. These definitions also hold when $u,v$ and $w$ are not distinct vertices.
If a solution $(z_{S})$ satisfies the 3-round Sherali-Adams constraints, then we would have
$y_{uv}+y_{uw}+y_{vw}=1+(2y_{uvw}-y_{u|v|w})$
and $y_{u|v|w}\in[0,1]$. However, for a solution to cluster LP, we have the following, weaker, claim.
Lemma 30.
The following holds in a solution of the cluster LP. For any $\{u,v,w\}$, we have
$3y_{uvw}\leq y_{uv}+y_{uw}+y_{vw}\leq\frac{3}{2}+\frac{3}{2}y_{uvw}.$
Proof.
For a vertex $u$, we have: $1-y_{uv}-y_{uw}+y_{uvw}\in[0,1]$. This follows from the fact that
$\sum_{S\ni u}z_{S}\geq\sum_{S\ni u,v}z_{S}+\sum_{S\ni u,w}z_{S}-\sum_{S\ni u,v,w}z_{S}\quad\iff\quad 1\geq y_{uv}+y_{uw}-y_{uvw}.$
Now we consider three equalities, one each for $u,v$ and $w$. Adding up and dividing by 3 implies the lemma. ∎
@@
$\Pr[\{v,w\}\subseteq C\mid u]$ $=\begin{cases}y_{uvw}&\text{if both $uv$ and $uw$ are long $+$edges}\\ \Pr[v\in C\mid u]\cdot\Pr[w\in C\mid u]&\text{otherwise}\end{cases},$
$\Pr[v\in C\wedge w\notin C\mid u]$ $=\begin{cases}y_{uv|w}&\text{if both $uv$ and $uw$ are long $+$edges}\\ \Pr[v\in C\mid u]\cdot\big(1-\Pr[w\in C\mid u]\big)&\text{otherwise}\end{cases}.$
We now prove some useful facts about the budget functions initially defined in Section 2.2.
$\mathrm{budget}^{+}_{\alpha}(x)=\frac{\alpha}{1-\alpha/2}\cdot\frac{x^{2}}{1+x},\qquad\text{and}\qquad\mathrm{budget}^{-}_{\alpha}(x)=\frac{\alpha}{1-\alpha/2}\cdot\frac{(1+2x)(1-x)}{2(1+x)},\qquad\forall x\in[0,1].$
@@
For any $\alpha\in[1,2)$, $\mathrm{budget}^{+}_{\alpha}(x)$ is convex and increasing. Moreover, $\mathrm{budget}^{-}_{\alpha}(x)$ is decreasing.
@@
Consider the first and second derivative of $\mathrm{budget}^{+}_{\alpha}(x)$,
$\frac{d}{dx}\mathrm{budget}^{+}_{\alpha}(x)=\frac{\alpha}{1-\alpha/2}\cdot\frac{x(2+x)}{(1+x)^{2}}\geq 0,$
$\frac{d^{2}}{d^{2}x}\mathrm{budget}^{+}_{\alpha}(x)=\frac{\alpha}{1-\alpha/2}\cdot\frac{2}{(x+1)^{3}}\geq 0.$
Here, we used that $\frac{\alpha}{1-\alpha/2}\geq 0$ and $x\geq 0$. Similarly, for $\mathrm{budget}^{-}_{\alpha}(x)$,
$\frac{d}{dx}\mathrm{budget}^{-}_{\alpha}(x)=\frac{\alpha}{1-\alpha/2}\cdot\frac{-x(x+2)}{(1+x)^{2}}\leq 0.$
We will use the following claim several times.
Claim 34.
For $x\geq\tau$, $\mathrm{budget}^{+}_{1.56}(x)\geq 2x.$
Proof..
Our goal is to show that
$\frac{2\alpha}{2-\alpha}\frac{x^{2}}{1+x}\geq 2x.$
Simplifying, we have
$\frac{x}{1+x}\geq\frac{2-\alpha}{\alpha}.$
Solving this inequality, we have
$x\geq\frac{2-\alpha}{2\alpha-2}.$
When $\alpha=1.56$, the inequality holds when $x\geq 0.392857$. $\Diamond$
We will prove the lemma by considering each possible triangle type separately. That is, degenerate triangles, triangles with no ++edge (-\!-\!- triangles), triangles with one ++edge (++\!-\!- triangles), triangles with two ++edges (+++\!+\!- triangles) and lastly, triangles with three ++edges (++++\!+\!+ triangles). Moreover, recall that the pivot rounding algorithm treats short ++edges differently than long ++edges. Therefore, for each type of triangle, we consider different cases that specify how many short and long ++edges the triangle contains.
Degenerate triangles.
@@
Fix budget functions $\mathrm{budget}^{+}\equiv\mathrm{budget}^{+}_{\alpha}$ and $\mathrm{budget}^{-}\equiv\mathrm{budget}^{-}_{\alpha}$, where $\alpha\geq 4/3$. For any degenerate triangle $T$, we have $\mathrm{cost}(T)\leq\Delta(T)$.
@@
For a -edge with value $x$ this is equivalent to
@@
This holds true since $\frac{\alpha(1+2x)}{(2-\alpha)(1+x)}\geq\frac{\alpha}{(2-\alpha)}\geq 1$, for any $\alpha\geq 1$. For short ++edges the cost is 0 and the inequality holds trivially. For long ++edges, with value $x>0.4$, we need,
@@
This holds true since $\frac{2\alpha x}{(2-\alpha)(1+x)}\geq\frac{2\alpha/3}{(2-\alpha)(1+1/3)}\geq 1$ for any $\alpha\geq 4/3$. ∎
5.1 ++++\!+\!+ Triangles
@@
Fix the budget function $\mathrm{budget}^{+}\equiv\mathrm{budget}^{+}_{\alpha}$, where $\alpha=1.56$. For any ++++\!+\!+ triangle $T$, we have $\mathrm{cost}(T)\leq\Delta(T)$.
@@
We distinguish between the following four cases: the triangle has three short edges, two short edges, one short edge or no short edge.
($s,$s)
@@
For a ++++\!+\!+ triangle $T=(u,v,w)$, assume that $uv$ is long and $uw$, vwvw are short. Let $a=x_{uv},b=x_{uw}$ and $c=x_{vw}$. Then
$\mathrm{cost}(T)$ $=2a,$
$\Delta(T)$ $=\mathrm{budget}^{+}(a)+\mathrm{budget}^{+}(b)+\mathrm{budget}^{+}(c)\geq\mathrm{budget}^{+}(a)\geq 2a.$
Here, we used that $\mathrm{budget}^{+}(a)\geq 2a$ by Claim 34.
@@
For a ++++\!+\!+ triangle $T=(u,v,w)$, assume that $uv$ is short and $uw$ and $vw$ are long, and again let $a=x_{uv},b=x_{uw}$ and $c=x_{vw}$.
$\mathrm{cost}(T)$ $=b+c+y_{vw}+y_{uw}-2\cdot y_{uvw}=2-2\cdot y_{uvw}$
$\Delta(T)$ $=\mathrm{budget}^{+}(b)+\mathrm{budget}^{+}(c)+\mathrm{budget}^{+}(a)\cdot(y_{vw}+y_{uw}-y_{uvw})$
$\geq\mathrm{budget}^{+}(b)+\mathrm{budget}^{+}(c).$
Because $b>\tau$, we have $\mathrm{budget}^{+}(b)\geq 2b$, by Claim 34. Hence,
$\Delta(T)-\mathrm{cost}(T)$ $\geq 2\cdot(b+c)-2+2\cdot y_{uvw}$
$=2\cdot(2-y_{uw}-y_{vw}+y_{uvw})-2$
@@
In the last inequality, we used that $y_{uw}+y_{vw}-y_{uvw}\leq 1$
@@
In this case, we use correlated rounding. For a ++++\!+\!+ triangle $T=(u,v,w)$, let $a=x_{uv},b=x_{uw}$ and $c=x_{vw}$. We have
$\mathrm{cost}(T)$ $=2(y_{uv}+y_{uw}+y_{vw})-6y_{uvw},$
$\Delta(T)$ $=\mathrm{budget}^{+}(a)(y_{uw}+y_{vw}-y_{uvw})+\mathrm{budget}^{+}(b)(y_{uv}+y_{vw}-y_{uvw})+\mathrm{budget}^{+}(c)(y_{uv}+y_{uw}-y_{uvw}).$
Let us fix $s=y_{uv}+y_{uw}+y_{vw}$ (which implies a+b+c=3sa+b+c=3-s), and let us also fix $y_{uvw}$. Then we have
$\Delta(T)=$ $\mathrm{budget}^{+}(a)(s-1+a-y_{uvw})+\mathrm{budget}^{+}(b)(s-1+b-y_{uvw})+\mathrm{budget}^{+}(c)(s-1+c-y_{uvw}).$
Since $\Delta(T)$ is convex in terms of $a,b$ and $c$, then subject to the constraint a+b+c=3sa+b+c=3-s, the function $\Delta(T)$ is minimized when $a=b=c=\frac{3-s}{3}$. Let us set $p=1-\frac{s}{3}$. We have
$\Delta(T)$ $\geq 3\cdot\mathrm{budget}^{+}(p)(s-1+p-y_{uvw})=3\cdot\mathrm{budget}^{+}(p)(\frac{2s}{3}-y_{uvw})=2s\cdot\mathrm{budget}^{+}(p)-3y_{uvw}\cdot\mathrm{budget}^{+}(p).$
Now we want to show that $\Delta(T)-\mathrm{cost}(T)\geq 0$.
$\Delta(T)-\mathrm{cost}(T)$ $\geq 2s\cdot\mathrm{budget}^{+}(p)-3y_{uvw}\cdot\mathrm{budget}^{+}(p)-2s+6y_{uvw}.$
So we want to show the following inequality.
$2s\cdot\mathrm{budget}^{+}(p)+6y_{uvw}$ $\geq 2s+3y_{uvw}\cdot\mathrm{budget}^{+}(p).$ (21)
Since all edges are long, we have $x_{uv}+x_{uw}+x_{vw}\geq 3\tau$, which implies that $s\leq 3-3\tau=1.8$, since $\tau=0.4$. Hence, $s\in[0,1.8]$ and $p\in[0.4,1]$, so $\mathrm{budget}^{+}(p)$ is in the range $\mathrm{budget}^{+}(\tau),\mathrm{budget}^{+}(1)]$. We will consider three cases, according to the value of $\mathrm{budget}^{+}(p)$. The first case is when $\mathrm{budget}^{+}(p)\in[1,2]$, in which case Inequality (21) clearly holds.
The second case is when $\mathrm{budget}^{+}(p)\in(2,\mathrm{budget}^{+}(1)]$. For ease of notation, let $\bar{b}=\mathrm{budget}^{+}(p)$ and assume $\bar{b}>2$. Rearranging, we derive the next inequalities, which are all equivalent.
$2s\cdot\bar{b}+6y_{uvw}$ $\geq 2s+3y_{uvw}\cdot\bar{b}$
$2s+2s(\bar{b}-2)$ $\geq 3y_{uvw}\cdot\bar{b}-6y_{uvw}$
$2s+2s(\bar{b}-2)$ $\geq 3y_{uvw}(\bar{b}-2).$
Since $s=y_{uv}+y_{uw}+y_{vw}\geq 3y_{uvw}$, the inequality holds.
The third and last case is when $\mathrm{budget}^{+}(p)\in[\mathrm{budget}^{+}(\tau),1)$. Again, for simplicity, let $\bar{b}=\mathrm{budget}^{+}(p)$. Then our goal is to prove the following for $\bar{b}\in[\mathrm{budget}^{+}(\tau),1)$.
$2s\cdot\bar{b}+6y_{uvw}$ $\geq 2s+3y_{uvw}\cdot\bar{b}.$
Rearranging, we have
$3y_{uvw}+3y_{uvw}(1-\bar{b})$ $\geq 2s(1-\bar{b}).$
Again, rearranging, we have
$2s\leq 3y_{uvw}+\frac{3y_{uvw}}{1-\bar{b}}.$ (22)
By Lemma 30, we have $2s\leq 3+3y_{uvw}$. Thus, if we can show
$\frac{y_{uvw}}{1-\bar{b}}\geq 1\quad\iff\quad y_{uvw}\geq 1-\bar{b}$
then we have proved (22). To prove this, we will show
$1-\bar{b}\leq\frac{2s}{3}-1\leq y_{uvw}.$
The second inequality follows from Lemma 30. Recalling that $p=1-\frac{s}{3}$, we just need to show
$1-\bar{b}=1-\mathrm{budget}^{+}(1-\frac{s}{3})\leq\frac{2s}{3}-1,$
which is established next.
We want to show
$1-\mathrm{budget}^{+}_{1.56}(1-\frac{s}{3})\leq\frac{2s}{3}-1.$
Rearranging, we have
$2-\frac{2s}{3}\leq\mathrm{budget}^{+}_{1.56}(1-\frac{s}{3}),$
which follows from Claim 34, since $1-\frac{s}{3}\in[\tau,1]$. ∎
5.2 -\!-\!- Triangles
@@
Fix the budget function $\mathrm{budget}^{-}\equiv\mathrm{budget}^{-}_{\alpha}$, where $\alpha\geq 1$. For any -\!-\!- triangle $T$, we have $\mathrm{cost}(T)\leq\Delta(T)$.
@@
We have that $\mathrm{budget}^{-}_{\alpha}(x)=\frac{\alpha}{1-\alpha/2}\cdot\frac{(1+2x)(1-x)}{2(1+x)}\geq(1-x)$. As before, in the proof of Lemma 35, we have $\frac{\alpha(1+2x)}{(2-\alpha)(1+x)}\geq\frac{\alpha}{(2-\alpha)}\geq 1$, for any $\alpha\geq 1$. The inequality holds true if the coefficient in the budget is $1$.
@@
Fix budget functions $\mathrm{budget}^{+}\equiv\mathrm{budget}^{+}_{\alpha}$ and $\mathrm{budget}^{-}\equiv\mathrm{budget}^{-}_{\alpha}$, where $\alpha\geq\frac{4}{3}$. For any ++\!-\!- triangle $T$, we have $\mathrm{cost}(T)\leq\Delta(T)$.
@@
We will prove the following two cases.
short ++edge
For a ++\!-\!- triangle $T=(u,v,w)$, assume that $uv$ is a short ++edge and $uw$ and $vw$ are -edges. The following inequality holds when $\alpha\geq\frac{4}{3}$.
$\mathrm{budget}^{-}_{\alpha}(x)=\frac{\alpha}{1-\alpha/2}\cdot\frac{(1+2x)(1-x)}{2(1+x)}\geq 2(1-x).$
@@
$\Delta(T)$ $\geq\mathrm{budget}^{-}(b)+\mathrm{budget}^{-}(c)\geq 2(y_{uw}+y_{vw}).$
long ++edge
For a ++\!-\!- triangle $T=(u,v,w)$, assume that $uv$ is a long ++edge and $uw$ and $vw$ are -edges. For $x\geq\tau$, as observed in the proof of Lemma 35, we can lower bound the budget of the ++edge as follows.
$\mathrm{budget}^{+}_{\alpha}(x)=\frac{\alpha}{1-\alpha/2}\cdot\frac{x^{2}}{1+x}\geq\frac{2\alpha\tau}{(2-\alpha)(1+\tau)}x\geq x.$
As in the previous case, $\mathrm{budget}^{-}(x)\geq 2(1-x)$. The inequality $\mathrm{cost}(T)\leq\Delta(T)$ is true if the coefficient for the ++edge is $1$ and for the -edge is $2$.
$\mathrm{cost}(T)$ $=y_{uw}y_{uv}+y_{vw}y_{uv}+y_{uw}+y_{vw}-2y_{uw}y_{vw}=(2-a)(y_{uw}+y_{vw})-2y_{uw}y_{vw}.$
$\Delta(T)$ $\geq(y_{uw}+y_{vw}-y_{uw}y_{vw})a+2(y_{uw}+y_{uv}-y_{uw}y_{uv})y_{vw}+2(y_{vw}+y_{uv}-y_{vw}y_{uv})y_{uw}$
$=(y_{uw}+y_{vw}-y_{uw}y_{vw})a+2(1-a+y_{uw}a)y_{vw}+2(1-a+y_{vw}a)y_{uw}$
$=(2-a)(y_{uw}+y_{vw})+3y_{uw}y_{vw}a.$
5.4 +++\!+\!- Triangles